You guy's don't have anything on adding integers (advanced). I don't understand it at all. && I just wanted to see if any of the tutor's have any strategies on solveing the problem's. The problems are things like 17 - (-23). Just simple'ishh thigs like that. Thanks for reading this. :)
Answer by CharlesG2(834)

You guy's don't have anything on adding integers (advanced). I don't understand it at all. && I just wanted to see if any of the tutor's have any strategies on solveing the problem's. The problems are things like 17 - (-23). Just simple'ishh thigs like that. Thanks for reading this. :)
17 - 23 = -6, why?
23 is 6 greater than 17
17 - (-23) = ?
now here we got 2 negative signs
so it is actually 17 + 23
17 + 23 = 40
2 negatives make a positive
if you subtract a larger number from a positive number your answer will be a negative number
if you add a larger number to a positive number your answer will be a positive number
if you subtract a negative number from a positive number your answer will be a positive number